OpenStudy (geometry_hater):

Find the area of the base of a pyramid with V = 360 and h = 12.

OpenStudy (anonymous):

90

OpenStudy (geometry_hater):

how

OpenStudy (anonymous):

360=base*12*1/3= 360=base*4 90=base

OpenStudy (anonymous):

V = (1/3)Bh Plug and solve.

OpenStudy (anonymous):

B = area of base

OpenStudy (geometry_hater):

thank you